OpenStudy (anonymous):

-7+n/22=-1 Show steps? answers:-15 (its hard for me to show my work; cause i do it in my head.)

OpenStudy (anonymous):

n/22 = 6 n = 6*22 = 132

OpenStudy (anonymous):

no -15 is the answer

OpenStudy (anonymous):

answer is not -15

OpenStudy (anonymous):

it's not right

OpenStudy (anonymous):

correct me if im wrong.

OpenStudy (anonymous):

the answer should be 132 not -15

OpenStudy (anonymous):

thats not even an option on my paper.

OpenStudy (anonymous):

So the problem is wrong

OpenStudy (mertsj):

\[\frac{-7+n}{22}\times22=-1\times22\]

OpenStudy (anonymous):

@smr_kml You're reading the question wrong. The question is as Mertsj has written it. The answer is correct.

OpenStudy (mertsj):

\[-7+n=-22\]

OpenStudy (anonymous):

The steps are as follows: Multiply both sides by 22, then add seven to both sides.

OpenStudy (anonymous):

\[-7+\frac{n}{22}=-1\] n/22 = -1 +7 n/22 = 6 n = 132

OpenStudy (mertsj):

\[-7+n+7=-22+7\]

OpenStudy (mertsj):

n=-15

OpenStudy (anonymous):

The takeaway lesson from this, @staceyyyy1, is to use parentheses.

OpenStudy (anonymous):

I got the same answer @JayDS

OpenStudy (anonymous):

Of course you got the same answer, smr, you're both solving the same incorrect problem.

OpenStudy (anonymous):

she has to put ( )

OpenStudy (anonymous):

Ok I'm sorry I didn't know that it's seperate

OpenStudy (mertsj):

Why doesn't the asker clarify??? We're more interested than the asker.

OpenStudy (mertsj):

@smr_kml There is A RAT in sepARATe

OpenStudy (anonymous):

To me it was plenty clarified by the fact that she said the answer was -15. However, she should in the future use parentheses for clarity.

OpenStudy (anonymous):

right! I'm sorry again everyone!

OpenStudy (anonymous):

thanks everyone. and thanks especially to Mertsj and nbouscal(: